# Cartan–Thullen theorem and Levi problem in context of generalised convexity

### v2

@@
Abstract.
We demonstrate that the Cartan—Thullen theorem and its generalisation to the context of generalised convexity, which we establish herein, can be regarded as a consequence of the classical theorems of functional analysis. Furthermore, we characterise the domains of holomorphy, and their generalisations, as the spaces that are complete, or as the spaces exhaustible by suitably defined polytopes. We also provide an analogue of the Levi problem and its elementary solution. Our result allows for a novel characterisation of Stein spaces.
Key words and phrases:
@@ Abstract.
2020 Mathematics Subject Classification:
Primary: 32E10, 32U05, 46N10, 52A01; Secondary: 31C05, 32E99, 32T35, 52B99
I would like to express my thanks to Robert McCann for bringing to my attention references regarding generalised convexity. I thank also Thomas Bloom for comments that helped improving the manuscript. Part of the research presented in this paper was performed when the author was a post-doctoral fellow at the University of Oxford, supported by the ERC Starting Grant CURVATURE, grant agreement No. 802689.
1. Introduction
The first aim of this article to show that the Cartan–Thullen theorem is an instance of a theorem in abstract generalised convexity, whose proof relies merely on functional analytic tools, such as the Banach-Alaoglu theorem and the Banach–Steinhaus uniform boundedness principle. The second aim is to provide a sensible formulation of the Levi problem in this setting, so that it covers several known results, and to resolve it. The resolution is elementary.
1.1. The Cartan–Thullen theorem and the Levi problem
@@
Let $\Omega\subset\mathbb{C}^{n}$ be an open set. For $B\subset\Omega$ its closed holomorphically convex hull is equal to
@@
The Cartan–Thullen theorem reads as follows, see [2] for the original paper and [21, Theorem 2.10.3 and Theorem 2.10.4, p. 77], [22, Theorem 3.4.5., p. 153] for more recent expoositions.
@@
      $\Omega_{2}$ is connected and $\Omega\setminus\Omega_{2}\neq\emptyset$,
@@
The Levi problem, posed by Levi in [24], was concerned with the question: is an open set $\Omega\subset\mathbb{C}^{n}$ such that any of its boundary points admits a neighbourhood $U$ such that $U\cap\Omega$ is holomorphically convex, necessarily holomorphically convex itself? The positive resolution of the problem is equivalent to demonstrating that holomorphic convexity of $\Omega$ is equivalent to the condition:
@@
This problem was solved by Oka [28] by homological methods, and independently by Bremermann [1] and Norguet [26]. Another approach, developed by Hörmander [18], [22, Chapter 5] employs methods of partial differential equations. We shall provide a simple proof of an analogue of this equivalence in the context of generalised convexity. We shall also prove a generalisation of Theorem 1.2.
We refer the reader also to a survey on the Levi problem [34], where also the Cartan–Thullen theorem is presented [34, Theorem 1.3., p. 482]. We refer to the book [22, Theorem 3.3.5., p. 144, Theorem 3.4.5., p. 153] for a thorough discussion of the problem in the setting of complex analysis. We refer also to [32] for a more recent paper on the topic as well as to [15] and [16, Chapter IX].
1.2. Generalised convexity
@@
We observe that not only does the notion of generalised convexity generalise classical convexity, when $\mathcal{F}$ is the cone of lower semi-continuous and convex functions, but also it includes the notion holomorphic convexity, when $\mathcal{F}$ is chosen to be the space of real-parts of holomorphic functions.
We refer the reader to the book of Hörmander [20] for an account comprising matter on notions of convexity with respect to: subharmonic functions, [20, Chapter III], plurisubharmonic functions, [20, Chapter IV]. The latter is equivalent to the notion of holomorphic convexity of complex analysis, cf. [22, Chapter 3].
@@
Let us mention that the notion of convexity has been developed and further generalised over the years, see [10], [30] and [33]. These developments mainly were concerned with various dualities, notions of subdifferentials and Fenchel transforms, which are parallel to the current developments.
1.2.1. Examples
In the book [20] Hörmander exhibits several examples that show that the equivalence of (i) of Theorem 1.2 and the Levi-type condition (iv) holds true also in a variety of settings. Let us list these equivalences:
@@ 1.2.1. Examples
    [20, Corollary 2.1.26., p. 59] shows that for an open set $U\subset\mathbb{R}^{n}$ its convexity is equivalent to the existence of a convex, proper function $p\colon U\to\mathbb{R}$,
@@
    [20, Theorem 3.2.31., p. 165] shows that if $U\subset\mathbb{R}^{n}$ is an open set, then the following conditions are equivalent:
@@
    [20, Theorem 4.1.19., p. 236] shows that if $U\subset\mathbb{C}^{n}$ is an open set, then the following conditions are equivalent:
@@
We shall see that these results of Hörmander are instances of general Theorem 1.18, in which we consider cones $\mathcal{F}$ of functions that satisfy the maximum principle or that are complete lattices generated by real-parts of a complex algebra.
We shall also see that the Cartan–Thullen theorem, Theorem 1.2, admits a generalisation to these settings.
1.3. Main results
Below we provide definitions of notions that are analogous to the notions of domain of holomorphy and domain of existence. These are Definition 1.5, (ii), and Definition 1.6 respectively.
@@ 1.3. Main results
The above defined topology $\tau(\mathcal{A})$ is generated by the basis sets of the form
@@
    $\Omega$ is complete with respect to $\mathcal{A}$ whenever every Cauchy net with respect to $\mathcal{A}$ in $\Omega$ is convergent in $\tau(\mathcal{A})$.
@@
Let $\Omega$ be a topological space and let $\mathcal{A}$ be a linear space of continuous functions on $\Omega$. Suppose that the topology on $\Omega$ agrees with $\tau(\mathcal{A})$. Suppose that $\Omega$ is metrisable with metric $d$ such that the Cauchy sequences with respect to $\mathcal{A}$ and the Cauchy sequences with respect to $d$ coincide. Let us denote by $\bar{\Omega}$ the completion of $\Omega$, i.e., the space of equivalence classes of all Cauchy sequences of elements in $\Omega$, where two Cauchy sequences $(\alpha_{n})_{n=1}^{\infty}$ and $(\beta_{n})_{n=1}^{\infty}$ are said to be equivalent if and only if
$\lim_{n\to\infty}d(\alpha_{n},\beta_{n})=0.$
We shall say that $\Omega$ is an $\mathcal{A}$-space whenever there exists $a\in\mathcal{A}$ that does not extend to a continuous function on $\Omega\cup\{\omega\}$ for any element $\omega\in\bar{\Omega}\setminus\Omega$.
@@
Suppose that $a\in\mathcal{A}$, $\omega\in\bar{\Omega}$. If $(\omega_{n})_{n=1}^{\infty}\subset\Omega$ is a Cauchy sequence with respect to $\mathcal{A}$, in the equivalence class of $\omega$, then we may put
$\bar{a}(\omega)=\lim_{n\to\infty}a(\omega_{n}).$
It is readily verified that this definition does not depend on the choice of the Cauchy sequence in the equivalence class of $\omega$. Thus, any $a\in\mathcal{A}$, extends to a continuous function $\bar{a}$ on $\Omega\cup\{\omega\}$.
Thus, $\Omega$ is complete with respect to $\mathcal{A}$ if and only if for any $\omega\in\bar{\Omega}$, there exists $a\in\mathcal{A}$ that does not extend to $\omega$ continuously.
$\Omega$ is an $\mathcal{A}$-space if and only if there exists $a\in\mathcal{A}$, that works for any $\omega\in\bar{\Omega}$.
Clearly, any $\mathcal{A}$-space is complete with respect to $\mathcal{A}$. Conversely, if $\Omega$ is complete with respect to $\mathcal{A}$, then it is an $\mathcal{A}$-space, as in this case $\bar{\Omega}=\Omega$.
@@
Suppose that $\Omega\subset\mathbb{C}^{n}$ is an open set and that $\mathcal{A}$ is the space of real-parts of holomorphic functions on $\Omega$. If $\Omega$ is complete with respect to $\mathcal{A}$, then it is a domain of holomorphy. Indeed, suppose there exist open sets $\Omega_{1},\Omega_{2}\subset\mathbb{C}^{n}$ such that $\emptyset\neq\Omega_{1}\subset\Omega_{2}\cap\Omega$, $\Omega_{2}$ is connected, $\Omega\setminus\Omega_{2}\neq\emptyset$, with the property that any holomorphic $a$ on $\Omega$ extends to a holomorphic function on $\Omega_{2}$. Let $\omega\in\partial\Omega\cap\Omega_{2}$. Since any holomorphic $a$ on $\Omega$ extends to a continuous function on $\Omega_{2}$, any sequence $(\omega_{n})_{n=1}^{\infty}\subset\Omega_{2}$ converging to $\omega$, is a Cauchy sequence with respect to $\mathcal{A}$. However, it is not convergent in $\Omega$, so $\Omega$ is not complete.
Similarly, if $\Omega$ is an $\mathcal{A}$-space, then it is a domain of existence.
@@
Let $\mathcal{F}$ be a cone of functions on a set $\Omega$. We say that $P$ is an $\mathcal{F}$-polytope whenever there exists a finite sequence $(f_{i})_{i=1}^{k}$ of functions in $\mathcal{F}$ such that
@@
    $p=\sup\{\lvert a_{\alpha_{i}}\rvert\mid i=1,\dotsc,k\},$
@@
    there exists a non-negative, proper function $p$ on $\Omega$, bounded on compactae, such that
@@
If $\Omega$ is separable and $\mathcal{A}$ consists of real-parts of a complex algebra, then (i) is equivalent to $\Omega$ being an $\mathrm{cl}\mathcal{A}$-space.
@@
If $\Omega$ is $\sigma$-compact, locally compact, Hausdorff topological space and $\mathcal{A}$ consists of continuous functions that separate points of $\Omega$, then the topology of $\Omega$ is the weak topology $\tau(\mathcal{A})$ induced by $\mathcal{A}$ as we shall see in Lemma 3.1.
@@
If $\Omega$ is $\sigma$-compact, locally compact, Hausdorff topological space and $\mathcal{A}$ is a closed algebra of continuous functions that separates points of $\Omega$, then $\mathcal{A}$ is the space of all continuous functions on $\Omega$, as follows by the Stone–Weierstrass theorem. Therefore we do not consider in the theorem the analogues of the statements for real algebras.
@@
We can think of function $p$ of (ii), (iii) or of (iv) as an analogue of a norm on the space $\Omega$. If $\mathcal{A}$ is the space of continuous linear functions on a finite-dimensional linear space $\Omega$, then $p$ is a genuine norm on $\Omega$.
@@
Let us also observe that the modulus, that appears in (i), does play a rôle. Indeed, let us fix a compact set $K\subset\Omega$. Proposition 4.7 shows that for $C\geq 1$ the sets
@@
1.3.1. Families satisfying maximum principle

### v3

@@
Abstract
We demonstrate that the Cartan–Thullen theorem and its generalisation to the context of generalised convexity, which we establish herein, can be regarded as consequences of the classical theorems of functional analysis: the Banach–Steinhaus theorem and the Banach–Alaoglu theorem. Furthermore, we characterise the domains of holomorphy, and their generalisations, as the spaces that are complete, or as the spaces exhaustible by suitably defined polytopes. We also provide an abstract analogue of the Levi problem and its elementary resolution. Our results allow for a novel characterisation of Stein spaces as the holomorphically complete spaces, as well as a proof that the Bremermann–Lelong lemma is equivalent to the positive answer to the Levi problem. Another contribution of ours is the introduction of the analogues of the notions of the complex analysis to the setting of generalised convexity.
keywords:
@@ Abstract
pacs:
MSC Classification (2020)]Primary: 32A70, 52A01; Secondary: 32Exx, 32T05, 32T35, 32U05, 46N10
[…]
Acknowledgments I would like to express my thanks to Robert McCann for bringing to my attention references regarding generalised convexity. I thank also Thomas Bloom for the comments that helped improving the manuscript. Part of the research presented in this paper was performed when the author was a post-doctoral fellow at the University of Oxford, supported by the ERC Starting Grant CURVATURE, grant agreement No. 802689.
1 Introduction
Our first aim is to show that the Cartan–Thullen theorem is an instance of a novel theorem in abstract generalised convexity, whose proof relies solely on functional analytic tools, such as the Banach-Alaoglu theorem and the Banach–Steinhaus uniform boundedness principle.
The second aim is to formulate the Levi problem in this abstract setting, and to resolve it. Our resolution relies on elementary soft analytic arguments. The resolution implies several results presented in the book of Hörmander [23] as particular instances.
We therefore show that the analogies between classical convexity and holomorphic convexity, as exposed in [1] or in [33], pertain to a much broader class of notions of generalised convexity.
The approaches to the original Levi problem either used homological methods, due to Oka [31], Bremermann [3] and Norguet [29], or involved partial differential equations, due to Hörmander [21]. In our developments the exhaustion function, which is the matter of concern in the Levi problem, is directly constructed using the very definition of holomorphic convexity, or the analogue of such convexity in the generalised convexity setting. [^1] In the general abstract setting of a topological space $\Omega$ and a linear space of continuous functions $\mathcal{A}$ on $\Omega$, studied in this paper, there is no clear notion of functions in $\mathcal{A}$ on the open subsets of $\Omega$ – unless the restrictions are considered, in which case the Levi problem becomes trivial. Therefore we consider the formulation concerning the exhaustion functions as the most appropriate generalisation of the Levi problem to our abstract setting. However, we do not obtain a new resolution of the original Levi problem.
A further contribution presented in this paper is the introduction of the analogues of the notions of domain of holomorphy and domain of existence to the setting of generalised convexity. Also notions of completeness and exhaustibility by polytopes are provided and their equivalence to the completeness with respect to the considered class of functions is shown. In the complex analytic setting it is proven that the holomorphic completeness is equivalent to Steinness of a complex analytic space. This shows that the intention of Stein to name the Stein spaces as holomorphically complete, see [40], was accurate.
1.1 The Cartan–Thullen theorem and the Levi problem
@@
Let $\Omega\subset\mathbb{C}^{n}$ be an open set. For $B\subset\Omega$ its closed holomorphically convex hull is defined as
@@
The Cartan–Thullen theorem reads as follows, see [5] for the original paper and [24, Theorem 2.10.3 and Theorem 2.10.4, p. 77], [25, Theorem 3.4.5., p. 153] for more recent expositions.
@@
      $\Omega_{2}$ is connected and $\Omega_{2}\setminus\Omega\neq\emptyset$,
@@
The Levi problem, posed by Levi in 1911 [27], was concerned with the following question: is an open set $\Omega\subset\mathbb{C}^{n}$ such that any of its boundary points admit a neighbourhood $U$ such that $U\cap\Omega$ is holomorphically convex, necessarily holomorphically convex itself? The positive resolution of the problem is equivalent to demonstrating that the holomorphic convexity of $\Omega$ is equivalent to the following condition:
@@
The problem remained unresolved for more than forty years until an affirmative resolution by Oka in 1950 [31], who used homological methods. Independent solutions were provided by Bremermann [3] and by Norguet [29]. Another approach, developed by Hörmander [21], see also [25, Chapter 5], employed methods of partial differential equations. We shall provide a simple proof of an analogue of the equivalence of (i)-(iv) in the context of generalised convexity. In particular, we prove a generalisation of Theorem 1.2.
The reader is referred to a survey on the Levi problem [39], where the Cartan–Thullen theorem is also presented [39, Theorem 1.3., p. 482]. For a modern discussion of the problem we refer to [25, Theorem 3.3.5., p. 144, Theorem 3.4.5., p. 153]. A more recent paper on the topic is [37].
Using our abstract formulation of the Levi problem we show that the classical lemma of Bremerman and Lelong [4, Theorem 2] is equivalent to the affirmative answer to the Levi problem.
1.2 Generalised convexity
@@
Not only does the notion of generalised convexity generalise classical convexity, when $\mathcal{F}$ is the cone of lower semi-continuous and convex functions, but it also includes the notion of holomorphic convexity, when $\mathcal{F}$ is chosen to be the space of real-parts of holomorphic functions.
The reader is referred to the book of Hörmander [23] for an account comprising matter on notions of convexity with respect to: subharmonic functions, [23, Chapter III], plurisubharmonic functions, [23, Chapter IV]. The latter is equivalent to the notion of holomorphic convexity of complex analysis, cf. [25, Chapter 3].
@@
The notion of convexity has been developed and further generalised over the years, see [15], [35], and [38]. These developments were mainly concerned with various dualities, notions of subdifferentials and Fenchel transforms, which are parallel to the current developments.
1.2.1 Examples
In the book [23] Hörmander exhibits several examples that show that the equivalence of (i) of Theorem 1.2 and the Levi-type condition (iv) holds true in a variety of settings that involve generalised convexity. These equivalences are:
@@ 1.2.1 Examples
    [23, Corollary 2.1.26., p. 59] shows that for an open set $U\subset\mathbb{R}^{n}$ its convexity is equivalent to the existence of a convex, proper function $p\colon U\to\mathbb{R}$,
@@
    [23, Theorem 3.2.31., p. 165] shows that if $U\subset\mathbb{R}^{n}$ is an open set, then the following conditions are equivalent:
@@
    [23, Theorem 4.1.19., p. 236] shows that if $U\subset\mathbb{C}^{n}$ is an open set, then the following conditions are equivalent:
@@
We shall see that these results of Hörmander are instances of general Theorem 1.18, in which we consider cones $\mathcal{F}$ of functions that satisfy the maximum principle. Another related result is Theorem 1.11 concerned with complete lattices generated by a linear subspaces of functions, see Definition 2.7.
We mention also the work of Range [33] presenting relations between pseudoconvexity and convexity.
We shall also see that the Cartan–Thullen theorem, Theorem 1.2, admits a generalisation to the setting of generalised convexity, under the assumption that the cone $\mathcal{F}$ is generated by a linear space of functions, see Definition 2.7, that is, any function in the cone is the supremum of a collection of functions in the linear space.
1.3 Domains of holomorphy and domains of existence
Below novel definitions are provided serving as analogues of the notions of domain of holomorphy and domain of existence. These are Definition 1.5, (ii), and Definition 1.6 respectively.
@@ 1.3 Domains of holomorphy and domains of existence
The topology $\tau(\mathcal{A})$ defined above is generated by the basis sets of the form
@@
    $\Omega$ is complete with respect to $\mathcal{A}$ whenever every Cauchy net with respect to $\mathcal{A}$ in $\Omega$ is convergent in $\tau(\mathcal{A})$ to an element in $\Omega$.
@@
Let $\Omega$ be a topological space and let $\mathcal{A}$ be a linear space of functions on $\Omega$. Let $d$ be a metric on $\Omega$ that generates the same topology as $\tau(\mathcal{A})$.
We shall say that $\Omega$ is an $(\mathcal{A},d)$-space whenever there exists a single function $a\in\mathcal{A}$ for which
$\lim_{\alpha\in A}a(\omega_{\alpha})=\infty$
for any Cauchy net $(\omega_{\alpha})_{\alpha\in A}$ with respect to $d$ in $\Omega$ that does not converge in $\tau(\mathcal{A})$ to an element in $\Omega$.
@@
Note that it is not always true that a function $a\in\mathcal{A}$ will extend to a continuous function on the completion of $\Omega$. However, it is so, if, for example, we assume that $a$ is uniformly continuous.
@@
Consider an open set $\Omega\subset\mathbb{C}^{n}$ and take $\mathcal{A}$ to be the space of real-parts of holomorphic functions on $\Omega$. If $\Omega$ is complete with respect to $\mathcal{A}$, then it is a domain of holomorphy. Indeed, suppose there exist open sets $\Omega_{1},\Omega_{2}\subset\mathbb{C}^{n}$ such that $\emptyset\neq\Omega_{1}\subset\Omega_{2}\cap\Omega$, $\Omega_{2}$ is connected, $\Omega_{2}\setminus\Omega\neq\emptyset$, with the property that for any holomorphic $a$ on $\Omega$, its restriction to $\Omega_{1}\cap\Omega$ extends to a holomorphic function on $\Omega_{2}$. Let $\omega\in\partial\Omega\cap\Omega_{2}$. By the extension properties of the holomorphic functions on $\Omega$, any sequence $(\omega_{n})_{n=1}^{\infty}\subset\Omega_{2}$ converging to $\omega$, is a Cauchy sequence with respect to $\mathcal{A}$. However, it is not convergent in $\Omega$, so $\Omega$ is not complete with respect to $\mathcal{A}$.
Similarly, if $\Omega$ is an $(\mathcal{A},d)$-space, where $d$ is the standard Euclidean metric, then it is a domain of existence. Let $a\in\mathcal{A}$ be a function such that
$\lim_{\alpha\in A}a(\omega_{\alpha})=\infty$
for any Cauchy net $(\omega_{\alpha})_{\alpha\in A}$ with respect to $d$ in $\Omega$ that does not converge in $\tau(\mathcal{A})$. Then there are no open sets $\Omega_{1},\Omega_{2}\subset\mathbb{C}^{n}$ such that $\emptyset\neq\Omega_{1}\subset\Omega_{2}\cap\Omega$, $\Omega_{2}$ is connected, $\Omega_{2}\setminus\Omega\neq\emptyset$, with the property that the restriction of $a$ to $\Omega\cap\Omega_{1}$ extends to a holomorphic function on $\Omega_{2}$.
1.4 Main results
@@
Let $\mathcal{F}$ be a cone of functions on a set $\Omega$. We say that $P$ is an $\mathcal{F}$-polytope whenever there exists a finite family $(f_{i})_{i=1}^{k}$ of functions in $\mathcal{F}$ such that
@@
In the first of our main results, we provide an abstract version of the Levi problem, which is the equivalence of (i)-(iv) below. Furthermore, under the additional assumption of separability of $\Omega$, we prove an abstract generalisation of the Cartan–Thullen theorem, cf. Theorem 1.2: we show that conditions (i)-(iv) are equivalent to $\Omega$ being complete with respect to $\mathrm{cl}\mathcal{A}$, which is an analogous notion to the notion of domain of holomorphy. If $\mathcal{A}$ consists of real-parts of a complex algebra, then these conditions are also equivalent to $\Omega$ being an $(\mathrm{cl}\mathcal{A},d)$-space, for any $d$ that metrises $\Omega$, which is an analogue of the notion of domain of existence.
@@
    $p=\sup\{\lvert a_{\alpha_{i}}\rvert\mid i=1,\dotsc,k\}$
    on $K$,
@@
    there exists a non-negative, proper function $p$ on $\Omega$, bounded on compacta, such that
@@
If $\Omega$ is separable and $\mathcal{A}$ consists of real-parts of a complex algebra and $d$ is any metric on $\Omega$ that generates the same topology as $\tau(\mathcal{A})$ , then (i) is equivalent to $\Omega$ being a $(\mathrm{cl}\mathcal{A},d)$-space.
@@
If $\Omega$ is $\sigma$-compact, locally compact, Hausdorff topological space and $\mathcal{A}$ consists of continuous functions that separate points of $\Omega$, then the topology of $\Omega$ is the weak topology $\tau(\mathcal{A})$ induced by $\mathcal{A}$, as we shall see in Lemma 3.1.
@@
If $\Omega$ is $\sigma$-compact, locally compact, Hausdorff topological space and $\mathcal{A}$ is a closed algebra of real-valued continuous functions that separates points of $\Omega$, then $\mathcal{A}$ is the space of all continuous functions on $\Omega$, as follows by the Stone–Weierstrass theorem. Therefore we do not consider in the theorem the analogues of the statements for real algebras that were stated for complex algebras.
@@
The function $p$ of (ii), (iii) or of (iv) can be thought of as an analogue of a norm on the space $\Omega$. If $\mathcal{A}$ is the space of continuous linear functions on a finite-dimensional linear space $\Omega$, then $p$ is a genuine norm on $\Omega$.
@@
Let us, however, observe that the modulus, which appears in (i), does play a rôle in the general case. Indeed, let us fix a compact set $K\subset\Omega$. Proposition 4.10 shows that for $C\geq 1$ the sets
@@
1.4.1 Cones satisfying the maximum principle
The second main result is concerned with another abstract version of the Levi problem in the generalised convexity setting. Instead of considering linear spaces of functions, we are now investigating the case of convexity with respect to a convex cone of functions that satisfies the maximum principle.